OpenStudy (anonymous):

i need to find the interval of convergence of the series (from 1 to infinity) of ((-1^n+1)(x-2)^n)/n2^n and then check for the convergence at the endpoints. i got the interval but i am having trouble with checking the endpoints. help ):

OpenStudy (anonymous):

\[\sum_1^{\infty}\frac{(-1)^{n+1}(x-2)^n}{n2^n}\]

OpenStudy (anonymous):

i am going to make a guess that the radius is 2, so that you will have \[|x-2|<2\] is that what you got?

OpenStudy (anonymous):

YES. and then i added 2 to both sides. so 0<x<4 ?

OpenStudy (anonymous):

that is what i got but i kind of guessed. ok so now we have to check at x = 0 and x = 4 right?

OpenStudy (anonymous):

yerp....

OpenStudy (anonymous):

maybe i am off by a minus sign, not sure, but this is an alternating series so surely converges as the terms go to zero, so we can include 0

OpenStudy (anonymous):

wait. how did you do that. how does it cancel off?

OpenStudy (anonymous):

i cancled the \[2^n\]

OpenStudy (anonymous):

oooh lol sorry

OpenStudy (anonymous):

my mistake i meant \[\sum\frac{(-1)^n}{n}\] oops

OpenStudy (anonymous):

i canceled wrong...

OpenStudy (anonymous):

no actually my algebra is terrible, can you explain the cancelling please?

OpenStudy (anonymous):

but conclusion is right. converges because you have an alternating series wehre terms go to zero

OpenStudy (anonymous):

you have in the numerator \[(-2)^n\] and in the denominator you have \[2^n\] but \[(-2)^n=(-1)^n2^n\] so you can cancel a \[2^n\]top and bottom

OpenStudy (anonymous):

kick the minus sign in to the first part to make it \[(-1)^n\] instead of \[(-1)^{n+1}\]

OpenStudy (anonymous):

that just leaves you with \[\sum\frac{(-1)^n}{n}\]

OpenStudy (anonymous):

ohh okay. and then by alternating series test it converges. got it! what about at x=4?

OpenStudy (anonymous):

well i didn't do that yet, so lets try it

OpenStudy (anonymous):

haha okay.

OpenStudy (anonymous):

i think this one is a problem because it will not alternate

OpenStudy (anonymous):

i think it will alternate? and it diverges because 1/n is harmonic.

OpenStudy (anonymous):

thank you so much though! do you know how to differentiate and integrate the series ones too?

OpenStudy (anonymous):

wait wait!!

OpenStudy (anonymous):

hold on a second i think i might have made a big big mistake!!

OpenStudy (anonymous):

why? i think it is correct...

OpenStudy (anonymous):

let me be more careful with zero

OpenStudy (anonymous):

whoa did i make a mistake!! lets go slow

OpenStudy (anonymous):

oh oops i think the x=4 is convergent too.

OpenStudy (anonymous):

lets look at the numerator when x = 0

OpenStudy (anonymous):

\[(-1)^{n+1}(-2)^n\] \[(-1)^{n+1}(-1)^n2^n\] \[(-1)^{2n+1}2^n\]

OpenStudy (anonymous):

now we can certainly cancel a \[2^n\] with the denominator

OpenStudy (anonymous):

get \[\sum\frac{(-1)^{2n+1}}{n}\]

OpenStudy (anonymous):

ohh i seee...

OpenStudy (anonymous):

ratio test?

OpenStudy (anonymous):

but \[2n+1\] is odd, and so the numerator is -1

OpenStudy (anonymous):

in other words you get \[\sum\frac{-1}{n}\] which diverges for sure

OpenStudy (anonymous):

ohhh okay okay. sounds good. um can you help me with another one?

OpenStudy (anonymous):

sorry about that , it was a big mistake on my part. i just thought to kick over the minus sign, but i forgot to add the exponents

OpenStudy (anonymous):

we did not do it with 4 yet, wanna try it?

OpenStudy (anonymous):

4 is going to converge, because this time the numerator is \[(-1)^{n+1}2^n\]

OpenStudy (anonymous):

cancel the \[2^n\] get \[\sum \frac{(-1)^{n+1}}{n}\]

OpenStudy (anonymous):

and this one alternates, so it converges for x = 4 but not for x = 0

OpenStudy (anonymous):

oh okay i get it!!

OpenStudy (anonymous):

yeah well i made a mistake, sorry it took so long

OpenStudy (anonymous):

you were so much help regardless. i barely know what i am doing. can i ask you for help on another one?

OpenStudy (anonymous):

sure but post in a new thread and i will come look it is much easier than scrolling down

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

yw